OpenStudy (nightwielder):

Medal and fan- of course. I answered the question with correct answers. thanks @sangya21 During a research experiment it was found that the number of bacteria in a culture grew at a rate proportional to it's size. At 8:00 AM there were 2,000 bacteria present in the culture. At noon, the number of bacteria grew to 2,800. How many bacteria will there be by midnight? During a research experiment it was found that the number of bacteria in a culture grew at a rate proportional to it's size. At 10:00 AM there were 3,000 bacteria present in the culture. At noon, the number of bacteria grew to 3,500. How many bacteria will there be by midnight?

OpenStudy (nightwielder):

\[P(t)=P _{0}e ^{kt}\]

OpenStudy (nightwielder):

I could do it myself but I don't enjoy interest formulae and exponential growth problems.

OpenStudy (anonymous):

let time be at x and no. of bacteria be y slope of line = (2800 - 2000)/ (12-8) m = 800/4 m = 200 rate of growth is 200 bacteria per hour so from noon to midnight, there are 12 hours, 200 * 12= 2400 2400 more bacteria will be there

OpenStudy (nightwielder):

becuz 4 4 hrs it increases 800 so 800/4 right

OpenStudy (anonymous):

correct

OpenStudy (nightwielder):

ok. I'm sorry for being lazy.

OpenStudy (nightwielder):

I'll probalby have some more problems so I might tag you.

OpenStudy (anonymous):

Okie dokie

OpenStudy (anonymous):

just remember to close this question

OpenStudy (nightwielder):

2000+2400=4400 for final answer

OpenStudy (nightwielder):

of course. im not new to openstudy.

OpenStudy (nightwielder):

oops. 2800+2400=5200

OpenStudy (nightwielder):

ugh. corect answer is 7,683.

OpenStudy (nightwielder):

I pressed similar excersise for a second chance.

OpenStudy (nightwielder):

During a research experiment it was found that the number of bacteria in a culture grew at a rate proportional to it's size. At 10:00 AM there were 3,000 bacteria present in the culture. At noon, the number of bacteria grew to 3,500. How many bacteria will there be by midnight?

OpenStudy (nightwielder):

^^^new question

OpenStudy (anonymous):

2800+2400

OpenStudy (anonymous):

how will u approach this question?

OpenStudy (nightwielder):

using the equation: \[3500=3000(e ^{2k})\]

OpenStudy (nightwielder):

i skipped a whole bunch of steps`

OpenStudy (nightwielder):

im in College Algebra btw

OpenStudy (anonymous):

Okk. I actually dont remember this equation so I take long route

OpenStudy (nightwielder):

\[\ln\frac{ 7 }{ 6 }=\ln(e ^{2k})\]

OpenStudy (nightwielder):

by using camcellation property f logs \[\ln \frac{ 7 }{ 6 }=2k\]

OpenStudy (nightwielder):

*of

OpenStudy (nightwielder):

*cancellation

OpenStudy (anonymous):

good. :)

OpenStudy (nightwielder):

ln(7/6)=0.15415 I'm rounding here but I won't round in the actual calculations.

OpenStudy (nightwielder):

then u divide by 2 k=0.077

OpenStudy (nightwielder):

so from 10:00 to noon is 2 hours. noon to midnight is 12. so total time is 14.`

OpenStudy (nightwielder):

the ` symbol is from me trying to press the tab for tab-enter

OpenStudy (anonymous):

so your total no. of bacteria will be????

OpenStudy (nightwielder):

\[P(14)=3,000(e ^{14k})\]

OpenStudy (nightwielder):

plugging in k.

OpenStudy (nightwielder):

P(14)=3,000(2.941897)

OpenStudy (nightwielder):

P(14)=8826

OpenStudy (nightwielder):

btw this is for anyone who needs help.

OpenStudy (nightwielder):

thats why i solved it here

OpenStudy (nightwielder):

it is correct!

OpenStudy (nightwielder):

I have no idea why I wrote fan. ok, @sangya21 is fanned -_-^-^-_-^-^-_-^-^-_-^-^-_-^-^-_-^-^-_-^-^-_-^-^-_-^-^-_-^-^-_-^-^

OpenStudy (nightwielder):

^^^those are my mixed emotions.

OpenStudy (anonymous):

anytime and your welcome. :)

OpenStudy (nightwielder):

thanks sooo much forever